coseca=4 and cota=root3 k
find k pppppppppppppppppplllllllllllllllllllllllllllleeeeeeeeeeeeeaaaaaaaaaasssssssseeeeeeeee help
 

Grade:10

2 Answers

SAI SARDAR
8 years ago
IF COSEC A=4 AND COT A=ROOT3 * K THEN IF A IS EQUAL TO THE 300 THEN K VALUE IS 2.OTHER CONDITIONS ARE NOT SATISFIED.
Tejaswini
5 years ago
1+cot2theta=cosec2theta
1=cosec2theta-cot2theta
1=42-(root3×k)2    (substituting in the values
1=16-3k.                           Cosectheta=4 and  
3k=16-1.                      Cottheta=root3*k)
3k=15
k=15/3
k=5